# CoLLEGe: Concept Embedding Generation for Large Language Models

### v1

@@ Abstract
Current language models are unable to quickly learn new concepts on the fly, often requiring a more involved finetuning process to learn robustly. Prompting in-context is not robust to context distractions, and often fails to confer much information about the new concepts. Classic methods for few-shot word learning in NLP, relying on global word vectors, are less applicable to large language models. In this paper, we introduce a novel approach named CoLLEGe (Concept Learning with Language Embedding Generation) to modernize few-shot concept learning. CoLLEGe is a meta-learning framework capable of generating flexible embeddings for new concepts using a small number of example sentences or definitions. Our primary meta-learning objective is simply to facilitate a language model to make next word predictions in forthcoming sentences, making it compatible with language model pretraining. We design a series of tasks to test new concept learning in challenging real-world scenarios, including new word acquisition, definition inference, and verbal reasoning, and demonstrate that our method succeeds in each setting without task-specific training.
@@ 1 Introduction
Figure 1: Our model generates an embedding for an unseen token given two example sentences. The ground truth word is pendant, and the model is able to generate an accurate definition using the embedding produced by CoLLEGe.
Building intuitions about the meaning of unfamiliar concepts in this way, with only a few examples of their usage, is common in real-world human learning, but remains difficult for language models, particularly when we want to consolidate this knowledge into discrete tokens. Providing a few in-context examples of how to use these new tokens can be a stopgap, but is often less powerful, leads to increased context length, and the additional examples can at times serve as distractors that confuse the language model, to say nothing about how unnatural it is (imagine if, each time the professor wanted the student to answer a question about epistemological idealism, they began by repeating a few sentences containing “idealism”). Instead, with a few examples showing how a concept is used, language models should know general semantic knowledge about this new concept, similar to the knowledge encoded in their pretrained representations. We frame this as a few-shot learning problem, where, given a few example sentences, the goal is generate an embedding for a new concept token with expressive and task-general semantic information.
@@
Furthermore, prior evaluation methods for new concept learning were limited to noisy proxy measures, such as the correlation between embedding cosine similarity and human similarity judgements (Lazaridou et al., 2017), or the cosine similarity between the few-shot concept embeddings and “ground truth” Word2Vec embeddings (e.g. the definitional nonce in Herbelot & Baroni (2017)). Current language models are able to use language in highly sophisticated and complex ways; true evaluation of few-shot concept learning in this setting should assess whether new concepts can be used in similarly complex and sophisticated ways. What older evaluations measure (e.g. correlation with human similarities or with ground-truth embeddings) tell us little about how well language models can use learned concepts. How well can they define a concept given only a few examples? Can they correctly answer fill-in-the-blank questions for difficult words? We can, and often do, ask humans the same questions to determine how well they have internalized a new concept. Our evaluations of language models should follow suit.
@@
  • Challenging datasets to measure the effectiveness of few-shot concept learning methods for LLMs. They test both general and complex concept knowledge, naturalistic acquisition of new concepts, and relational abstraction;
  • Experiments showing that, by training an embedding generation modules in a task-general manner, we can generate embeddings that, without additional training, allow a pretrained LLM to: a) generate plausible definitions for new concepts, b) correctly solve fill-in-the-blank tasks for difficult words, and c) correctly identify the meaning of new slang terms.
Figure 2: Our proposed CoLLEGe framework for concept embedding generation. Support sequences are embedded by a pretrained masked language model (MLM) (e.g. RoBERTa) with an additional Transformer encoder to produce pooled sequence embeddings for each support sequence. These are aggregated and projected into the input and output embedding space for the pretrained LLM (e.g. LLaMA). According to UrbanDictionary, beige flag, an Internet slang appeared in mid-2023, means “a benign but annoying trait or habit.” [^2] https://www.urbandictionary.com/define.php?term=Beige%20flag
@@ 2 Related Work
A classic and related task in NLP is rare word learning (Luong et al., 2015). Lazaridou et al. (2017) create the synthetic “chimera” concepts and provide early evidence that summation over (global) word vectors in the context surrounding a new or rare word can produce a useful embedding. Khodak et al. (2018) build on this, presenting a method which includes a learned linear transformation to account for shared features across global word vectors. Lampinen & McClelland (2017) present an even simpler method, involving freezing the majority of the weights of the network and using gradient descent to tune only the weights related to the new word(s). For a more complex approach, Herbelot & Baroni (2017) modify the Word2Vec algorithm for more effective few-shot learning. More modern approaches include HiCE (Hu et al., 2019), which uses Transformer layers to induce a new word embedding from Word2Vec embeddings, and Mem2Vec (Sun et al., 2018), which uses a long-term memory system. Similarly, Weston et al. (2015) model new words by using contextual examples from memory. They store a bag-of-words for the left and right context surrounding the new word, and simulate new word learning with a fixed percent of words encountered during training. By contrast, we use a frozen MLM to learn a representation for the new concept token. Other approaches incorporate morphological information (Luong et al., 2013; Schick & Schütze, 2019). While these methods are useful, particlarly for learning global word vector representations, they are less useful for augmenting the embeddings of pretrained LLMs. In part, this is because global word vectors do not map easily to the pretrained LLM embedding space. Additionally, global word vector representations are often less informative than the pretrained representations from BERT-style models.
@@
Matching Networks (Vinyals et al., 2016) and Prototypical Networks (Snell et al., 2017) both approach few-shot learning as a meta-learning problem. Online prototypical networks (Ren et al., 2020) build on the latter for novel concept learning in an online and continual fashion. Our approach is also related to fast-weight networks (Schmidhuber, 1992), since we use the support sequences to generate a fast embedding weight for the new concept.
For language, meta-learning is often used for knowledge augmentation (Hu et al., 2023), task adaptation (Chen et al., 2022; Zhong et al., 2021; Bansal et al., 2020), domain adaptation (Qian & Yu, 2019; Li et al., 2020; Geng et al., 2019), rare word recognition (Lux & Vu, 2021) (in ASR), and word sense disambiguation Holla et al. (2020), among other applications (Lee et al., 2022). Some recent methods frame meta-learning as a sequence modeling problem, drawing inspiration from in-context learning (Chen et al., 2022; Fifty et al., 2023). Finally, Lake & Baroni (2023) recently developed a method for learning compositional concepts. In our work, context sentences are encoded to generate a new embedding, conceptually similar to a prototype for the new concept, to optimize a general language modeling objective, rather than a collection of task objectives.
@@
3 CoLLEGe: Concept Learning with Language Embedding
Generation
In this section, we describe our proposed approach for enabling LLMs to quickly learn new concept tokens. Given a new word or concept and a set of example sentences containing that word or concept, we want to produce an embedding that captures its semantically meaningful features.
@@ 3 CoLLEGe: Concept Learning with Language Embedding
To do this, the new token is replaced with a <mask> token in each support sequence, and each is embedded with a frozen masked language model (MLM) used for feature extraction. The contextual embeddings for each sequence are then passed through an additional Transformer self-attention layer to process the contextual embeddings for each sequence to obtain $\{\bm{h}_{i,t}\}$. These are then aggregated using mean pooling, producing $k$ sequence embeddings $\{\bm{e}_{1},...,\bm{e}_{k}\}$:
$\bm{e}_{i}=\frac{1}{n_{i}}\sum_{t=1}^{n_{i}}\bm{h}_{i,t},$ (1)
where $n_{i}$ is the length of each sequence. The sequence embeddings are aggregated once more using mean pooling, producing a single output embedding $\bm{e}_{\text{new}}$:
$\bm{e}_{\text{new}}=\frac{1}{K}\sum_{i=1}^{K}\bm{e}_{i}.$ (2)
Mean pooling can also facilitate incremental consolidation of new concepts without having to store past examples. To integrate the embedding with a frozen autoregressive LM, we apply two distinct linear layers to produce an input and output embedding for the new token $\bm{e}_{\text{in}}$ and $\bm{e}_{\text{out}}$:
$[\bm{e}_{\text{in}},\bm{e}_{\text{out}}]=\mathrm{Linear}(\bm{e}_{\text{new}}).$ (3)
The autoregressive LM’s input and output token embedding matrices are then expanded with these generated embeddings, and used to model the query sequence.
@@
We visualize this process for a simple language modeling example in Figure 2.
@@
To efficiently sample support and query sequences, we save sentences containing a new token in an example buffer to serve as support sequences, and when we encounter the same token being used again in the training corpus, we will use the sequence as a query sequence. The query sequence can then be saved in the example buffer and used as a support sequence again. We find that reusing query sequences as support sequences is helpful for training and allows the model to make use of examples it has already learning when learning new examples. Often, query sequences are longer, comprising a few sentences or a paragraph of text. The sentence which contains the new token is extracted from each and used as a support sequence for a different query sequence concerning the same concept. Not every such sequence ends up in the final set, however, since we filter and rank the examples, see Section 4 for details.
@@
Initial experiments training on only positive examples, i.e. examples containing the new token, tended to yield generated embeddings with abnormally high norm compared to other LLM input or output token embeddings. One hypothesis was that, since every example the model has to learn contains a new token, it converges to an abnormally high norm embedding, to ensure that the new token appears in the query sequence. During normal pretraining, few tokens are shared across all sequences, and language models learn both when to generate and when not to generate each token. To likewise teach our model when not to generate a new token, we sample a sequence without a new token, which we call a negative example, and take the sum of the cross entropy loss on the positive example, $L_{\text{ce}}^{+}$, and the cross entropy loss on the negative example, $L_{\text{ce}}^{-}$.
@@
Since the pretrained LLM has already seen all the words before, we know the “true” language model embeddings and logits for the remainder of the sequence. Ideally, we want the generated embeddings from our model to match the ground truth embeddings and logits as faithfully as possible, to better approximate the underlying language model distribution. To do this, we retain the original sequence, before masking a word with a new token, and compute the output embeddings and logits for the rest of the sequence with a non-augmented version of our pretrained autoregressive model. We then compute the cosine distance between those output embeddings and the output embeddings from CoLLEGe, $L_{\text{cos}}$, as well as the MSE between the CoLLEGe LLM logits and the “true” LLM logits using the original embeddings, $L_{\text{mse}}$. Using a more standard objective with a distillation temperature (Hinton et al., 2015) was less effective during training.
@@
${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@rgb@stroke{0}{0}{0}\pgfsys@color@rgb@fill{0}{0}{0}L_{\text{cos}}}$ ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@rgb@stroke{0}{0}{0}\pgfsys@color@rgb@fill{0}{0}{0}=\frac{1}{n-|I_{\text{new}}|}\sum_{k\not\in I_{\text{new}}}1-\cos(\bm{e}_{t_{k,+}},\bm{e}_{t_{\sigma(k),\text{orig}}})},$ (6)
${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@rgb@stroke{0}{0}{0}\pgfsys@color@rgb@fill{0}{0}{0}L_{\text{mse}}}$ ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@rgb@stroke{0}{0}{0}\pgfsys@color@rgb@fill{0}{0}{0}=\frac{1}{n-|I_{\text{new}}|}\sum_{k\not\in I_{\text{new}}}(\bm{\ell}_{t_{k,+}}-\bm{\ell}_{t_{\sigma(k),\text{orig}}})^{2}},$ (7)
@@
Our final training loss is simply a sum of these individual loss terms. We explored using a linear combination of the loss terms to weight each differently, but found it to be at best no more effective than an unweighted sum and at times worse. Our final loss, $L_{\text{total}}$, is:
$L_{\text{total}}$ $=\underbrace{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@rgb@stroke{0}{0}{0}\pgfsys@color@rgb@fill{0}{0}{0}L_{\text{ce}}^{+}}+{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@rgb@stroke{0}{0}{0}\pgfsys@color@rgb@fill{0}{0}{0}L_{\text{ce}}^{-}}}_{\text{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@rgb@stroke{0}{0}{0}\pgfsys@color@rgb@fill{0}{0}{0}Cross Entropy Losses}}}+\underbrace{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@rgb@stroke{0}{0}{0}\pgfsys@color@rgb@fill{0}{0}{0}L_{\text{cos}}}+{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@rgb@stroke{0}{0}{0}\pgfsys@color@rgb@fill{0}{0}{0}L_{\text{mse}}}}_{\text{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@rgb@stroke{0}{0}{0}\pgfsys@color@rgb@fill{0}{0}{0}Distillation Losses}}}.$ (8)

### v2

@@ Abstract
Current language models are unable to quickly learn new concepts on the fly, often requiring a more involved finetuning process to learn robustly. Prompting in-context is not robust to context distractions, and often fails to confer much information about the new concepts. Classic methods for few-shot word learning in NLP, relying on global word vectors, are less applicable to large language models. In this paper, we introduce a novel approach named CoLLEGe (Concept Learning with Language Embedding Generation) to modernize few-shot concept learning. CoLLEGe is a meta-learning framework capable of generating flexible embeddings for new concepts using a small number of example sentences or definitions. Our primary meta-learning objective is simply to facilitate a language model to make next word predictions in forthcoming sentences, making it compatible with language model pretraining. We design a series of tasks to test new concept learning in challenging real-world scenarios, including new word acquisition, definition inference, and verbal reasoning, and demonstrate that our method succeeds in each setting without task-specific training. Code and data for our project can be found at https://college-concept-learning.github.io/.
@@ 1 Introduction
Building intuitions about the meaning of unfamiliar concepts in this way, with only a few examples of their usage, is common in real-world human learning, but remains difficult for language models, particularly when we want to consolidate this knowledge into discrete tokens. Providing a few in-context examples of how to use these new tokens can be a stopgap, but is often less powerful the additional examples can serve as distractors that confuse the language model, to say nothing about how unnatural it is (imagine if, each time the professor wanted the student to answer a question about epistemological idealism, they began by repeating a few sentences containing “idealism”). Instead, with a few example sentences, language models should know general semantic knowledge about this new concept, similar to the knowledge encoded in their pretrained representations. We frame this as a few-shot learning problem, where, given a few example sentences, the goal is generate an embedding for a new concept token with expressive and task-general semantic information.
Figure 1: Our model generates an embedding for an unseen token given one or a few example sentences. The ground truth word is pendant, and the model is able to generate an accurate definition using the embedding produced by CoLLEGe.
@@
Furthermore, prior evaluation methods for new concept learning were limited to noisy proxy measures, such as the correlation between embedding cosine similarity and human similarity judgements (Lazaridou et al., 2017), or the cosine similarity between the few-shot concept embeddings and “ground truth” Word2Vec embeddings (e.g. the definitional nonce in Herbelot & Baroni (2017)). Current language models are able to use language in highly sophisticated and complex ways; true evaluation of few-shot concept learning in this setting should assess whether new concepts can be used in similarly complex and sophisticated ways. What older evaluations measure tell us little about how well language models can use learned concepts. How well can they define a concept given only a few examples? Can they correctly answer fill-in-the-blank questions for difficult words? We ask humans the same questions to determine how well they have internalized a new concept. Our evaluations of language models should follow suit.
@@
  • Three challenging datasets – CoLLEGe-GRE, CoLLEGe-DefGen, and CoLLEGe-Slang – used to measure the effectiveness of few-shot concept learning methods for LLMs. These datasets test both general and complex concept knowledge, naturalistic acquisition of new concepts, and relational abstraction;
  • Experiments showing that, by training an embedding generation modules in a task-general manner, we can generate embeddings that allow a pretrained LLM to: a) generate plausible definitions for new concepts, b) correctly solve fill-in-the-blank tasks for difficult words, and c) correctly identify the meaning of new slang terms without additional training.
Figure 2: Our proposed CoLLEGe framework for concept embedding generation. Support sequences are embedded by a pretrained MLM (e.g. RoBERTa) with an additional Transformer encoder to produce pooled sequence embeddings for each support sequence. These are aggregated and projected into the input and output embedding space for the pretrained LLM (e.g. LLaMA). According to UrbanDictionary, beige flag, an Internet slang appeared in mid-2023, means “a benign but annoying trait or habit.” [^2] https://www.urbandictionary.com/define.php?term=Beige%20flag
@@ 2 Related Work
A classic and related task in NLP is rare word learning (Luong et al., 2015). Lazaridou et al. (2017) create the synthetic “chimera” concepts and provide early evidence that summation over (global) word vectors in the context surrounding a new or rare word can produce a useful embedding. Khodak et al. (2018) build on this, presenting a method which includes a learned linear transformation to account for shared features across global word vectors. Lampinen & McClelland (2017) present an even simpler method, involving freezing the majority of the weights of the network and using gradient descent to tune only the weights related to the new word(s). For a more complex approach, Herbelot & Baroni (2017) modify the Word2Vec algorithm for more effective few-shot learning. More modern approaches include HiCE (Hu et al., 2019), which uses Transformer layers to induce a new word embedding from Word2Vec embeddings, and Mem2Vec (Sun et al., 2018), which uses a long-term memory system. Similarly, Weston et al. (2015) model new words by using contextual examples from memory. They store a bag-of-words for the left and right context surrounding the new word, and simulate new word learning with a fixed percent of words encountered during training. Instead of using global word vectors to represent the context, we use the contextual representations from a frozen MLM to learn a representation for the new concept token. We also combine autoregressive losses, cosine distance, and distillation loss to train our model, whereas older work either reused the Skip-gram objective or used cosine distance. Other approaches incorporate morphological information (Luong et al., 2013; Schick & Schütze, 2019). While these methods are useful, particlarly for learning global word vector representations, they are less useful for augmenting the embeddings of pretrained LLMs. In part, this is because global word vectors do not map easily to the pretrained LLM embedding space. Additionally, global word vector representations are often less informative than the pretrained representations from BERT-style models.
@@
Matching Networks (Vinyals et al., 2016) and Prototypical Networks (Snell et al., 2017) both approach few-shot learning as a meta-learning problem. Online prototypical networks (Ren et al., 2021) build on the latter for novel concept learning in an online and continual fashion. Our approach is also related to fast-weight networks (Schmidhuber, 1992), since we use the support sequences to generate a fast embedding weight for the new concept. For language, meta-learning is often used for knowledge augmentation (Hu et al., 2023), task adaptation (Chen et al., 2022; Zhong et al., 2021; Bansal et al., 2020), domain adaptation (Qian & Yu, 2019; Li et al., 2020; Geng et al., 2019), rare word recognition (Lux & Vu, 2021) (in ASR), and word sense disambiguation Holla et al. (2020), among other applications (Lee et al., 2022). Some recent methods frame meta-learning as a sequence modeling problem, drawing inspiration from in-context learning (Chen et al., 2022; Fifty et al., 2024). Finally, Lake & Baroni (2023) recently developed a method for learning compositional concepts. In our work, context sentences are encoded to generate a new embedding, conceptually similar to a prototype for the new concept, to optimize a general language modeling objective, rather than a collection of task objectives.
@@
3 CoLLEGe: Concept Learning with Language Embedding Generation
In this section, we describe our proposed approach for enabling LLMs to quickly learn new concept tokens. Given a new word or concept and a set of example sentences containing that word or concept, we want to produce an embedding that captures its semantically meaningful features. We visualize this process for a simple language modeling example in Figure 2.
@@ 3 CoLLEGe: Concept Learning with Language Embedding Generation
To perform this generation, the new token in each support sequence is replaced with a <mask> token, and each is embedded with a frozen masked language model (MLM) used for feature extraction. The contextual embeddings for each sequence are then passed through an additional Transformer self-attention layer to process the contextual embeddings for each sequence to obtain $\{\bm{h}_{i,t}\}$. These are then aggregated using mean pooling, producing $k$ sequence embeddings $\{\bm{e}_{1},...,\bm{e}_{k}\}$: $\bm{e}_{i}=\frac{1}{n_{i}}\sum_{t=1}^{n_{i}}\bm{h}_{i,t},$ where $n_{i}$ is the length of each sequence. The sequence embeddings are aggregated once more using mean pooling, producing a single output embedding $\bm{e}_{\text{new}}$: $\bm{e}_{\text{new}}=\frac{1}{K}\sum_{i=1}^{K}\bm{e}_{i}.$ Mean pooling can also facilitate incremental consolidation of new concepts without having to store past examples. To integrate the embedding with a frozen autoregressive LM, we apply two distinct linear layers to produce an input and output embedding for the new token $\bm{e}_{\text{in}}$ and $\bm{e}_{\text{out}}$: $[\bm{e}_{\text{in}},\bm{e}_{\text{out}}]=\mathrm{Linear}(\bm{e}_{\text{new}}).$ The autoregressive LM’s input and output token embedding matrices are then expanded with these generated embeddings, and used to model the query sequence.
@@
To efficiently sample support and query sequences, terms we borrow from the few-shot learning literature, we save sentences containing a new token in an example buffer to serve as support sequences, and when we encounter the same token being used again in the training corpus, we will use the sequence as a query sequence. The query sequence can then be saved in the example buffer and used as a support sequence again. We find that reusing query sequences as support sequences is helpful for training and allows the model to make use of examples it has already learning when learning new examples. Often, query sequences are longer, comprising a few sentences or a paragraph of text. The sentence which contains the new token is extracted from each and used as a support sequence for a different query sequence concerning the same concept. Not every such sequence ends up in the final set, however, since we filter and rank the examples, see Section 4 for details.
@@
Initial experiments training on only positive examples, i.e. examples containing the new token, tended to yield generated embeddings with norms of much higher magnitude than those of other LLM input or output token embeddings. One hypothesis was that, since every example the model has to learn contains a new token, it prefers to converge an embedding with high norm, to ensure that the new token appears in the query sequence. During normal pretraining, few tokens are shared across all sequences, and language models learn both when to generate and when not to generate each token. To likewise teach our model when not to generate a new token, we sample a sequence without a new token, which we call a negative example, and take the sum of the cross entropy loss on the positive example, $L_{\text{ce}}^{+}$, and the cross entropy loss on the negative example, $L_{\text{ce}}^{-}$.
@@
Since we pretrain by replacing existing words with new tokens, we can compute the “true” language model embeddings and logits for the remainder of the sequence. Ideally, we want the generated embeddings from our model to match the ground truth embeddings and logits as faithfully as possible, to better approximate the underlying language model distribution. To do this, we retain the original sequence, before masking a word with a new token, and compute the output embeddings and logits for the rest of the sequence with a non-augmented version of our pretrained autoregressive model. We then compute the cosine distance between those output embeddings and the output embeddings from CoLLEGe, $L_{\text{cos}}$, as well as the MSE between the CoLLEGe LLM logits and the “true” LLM logits using the original embeddings, $L_{\text{mse}}$. Using a more standard objective with a distillation temperature (Hinton et al., 2015) was less effective during training.
@@
${\color[rgb]{0,0,0}L_{\text{cos}}}$ ${\color[rgb]{0,0,0}=\frac{1}{n-|I_{\text{new}}|}\sum_{k\not\in I_{\text{new}}}1-\cos(\bm{e}_{t_{k,+}},\bm{e}_{t_{\sigma(k),\text{orig}}})},$ (3)
${\color[rgb]{0,0,0}L_{\text{mse}}}$ ${\color[rgb]{0,0,0}=\frac{1}{n-|I_{\text{new}}|}\sum_{k\not\in I_{\text{new}}}(\bm{\ell}_{t_{k,+}}-\bm{\ell}_{t_{\sigma(k),\text{orig}}})^{2}},$ (4)
@@
Our final training loss is simply a sum of these individual loss terms. We explored using a linear combination of the loss terms to weight each differently, but found it was not significantly better. Our final loss, $L_{\text{total}}$, is:
$L_{\text{total}}$ $=\underbrace{{\color[rgb]{0,0,0}L_{\text{ce}}^{+}}+{\color[rgb]{0,0,0}L_{\text{ce}}^{-}}}_{\text{{\color[rgb]{0,0,0}Cross Entropy Losses}}}+\underbrace{{\color[rgb]{0,0,0}L_{\text{cos}}}+{\color[rgb]{0,0,0}L_{\text{mse}}}}_{\text{{\color[rgb]{0,0,0}Distillation Losses}}}.$ (5)